# From underground natural gas to hydrogen storage in fractured reservoir rock : comparing relative permeabilities for hydrogen versus methane and nitrogen

### v1

@@ 1 Introduction
The hydrogen economy is expected to play a critical role in the global transition to cleaner energy and achieving net-zero carbon emissions. A key technology supporting large-scale energy storage is Underground Hydrogen Storage (UHS). UHS enables the storage of TWh-scale energy by injecting hydrogen into depleted gas reservoirs, saline aquifers, and existing gas storage facilities. The stored (green) hydrogen can be produced by electrolysis during periods of surplus electricity generation [1, 2, 3, 4, 5, 6]. As renewable energy sources like wind and solar continue to expand [7], UHS provides a potential solution to balance renewable energy supply and demand, enhancing energy security, and helping to decarbonize energy systems.
@@
Converting pre-existing natural gas storage sites to hydrogen storage potentially offers advantages to using depleted gas fields or saline aquifers that have not been previously deployed for storage purposes. This includes leveraging expensive site characterization campaigns, tried-and-tested reservoir and caprock behaviour under cyclical injection-retraction operations, and existing connections to gas treatment and pipeline infrastructure. Among operational and planned natural gas storage sites in Europe, 86% are depleted gas fields and 14% are aquifers, together capable of providing up to 750 TWh of hydrogen storage [8]. Conversion of these storage resources would be sufficient for a mid-range scenario of 2,500 TWh of annual hydrogen demand by 2050, at a 30% storage capacity. In this paper, we focus on a notable example of such a natural gas storage site under investigation for hydrogen conversion: the Loenhout natural gas storage facility in northern Belgium. Loenhout is uniquely well-connected in Western Europe as it is part of a pipeline network linking Belgium, France, The Netherlands, Germany, Luxemburg and the United Kingdom, including international shipping terminals in the North Sea ports of Zeebrugge and Dunkirk. If converted succesfully, it would provide 2.2 TWh of storage capacity. The site comprises a partially karstified limestone aquifer from Visean age, characterized by a fracture-dominated flow system with low primary porosity [9, 10, 11]. Its conversion is currently under investigation in the BE-Hystore project funded by the Belgian federal Energy Transition Fund and the site operator Fluxys Belgium.
@@
Repurposing an aquifer natural gas storage site for large-scale hydrogen injection requires evaluating whether the flow of the stored gas in the pores or fractures of the storage reservoir would be affected, as this may impact injectivity and operational efficiency [12, 13]. While significant progress has been made in understanding hydrogen flow in porous media, substantial knowledge gaps remain regarding its behavior in fractured rock. This is not only important for fractured storage reservoirs such as in Loenhout, but also to assess leakage through caprock if fractures are present or generated during storage [14, 15]. A primary uncertainty in the current scientific understanding lies in hydrogen’s relative permeability compared to methane’s. Relative permeability is a key parameter for upscaled multiphase flow modeling, which quantifies how the flow of gas is impacted by the presence of (saline) groundwater in the reservoir, and vice versa [16].
@@
Over the last two decades, several studies have shown that the relative permeability behaviour of fractured media differs significantly from that of porous materials due to its dependence on complex fluid dynamics [17]. Nevertheless, it is currently unclear how such complexities affect fracture relative permeabilities for hydrogen compared to other gases, particularly at elevated pressures and temperatures typical for UHS. Experiments with air-water in transparent fractures have revealed strong phase interference and reduced relative permeability at intermediate saturations in rough fractures [18, 19]. Relative permeability in such systems is highly sensitive to fracture surface roughness and aperture heterogeneity, influencing the meandering or twisting of flow paths [20, 21, 22, 23, 24]. Capillary-dominated flow, especially in low viscosity-ratio fluid systems such as hydrogen-brine and methane-brine, is characterized by fluid pathways that are intermittently connected, caused by snap-off events and Haines jumps [25]. State-of-the-art fracture relative permeability models aim to capture the greater phase interference, the absence of percolating pathways, and altered flow patterns compared to porous media [26, 27, 28, 29, 30]. The current need for understanding how the findings above translate to UHS is underscored by a lack of experimental studies on fractured rock. This is despite significant recent efforts on the corresponding behaviour of porous reservoir rocks [31, 32, 33, 34, 35, 36, 37, 38], which are difficult to translate directly to fracture systems. Comparative studies of two-phase flow behavior of relevant fractured samples under consistent reservoir conditions are thus critical to address these uncertainties.
@@
Assessing hydrogen flow in fractured aquifers such as Loenhout is essential to evaluate its feasibility as a storage medium in comparison to methane. Therefore, our study presents the first direct measurements of drainage relative permeability for hydrogen-brine, methane-brine, and nitrogen-brine systems in fractured limestone rock at subsurface pressure conditions of 10 MPa. The investigated rock used was collected from a depth of at the Loenhout aquifer. Methane was selected because it is a good model for the stored gas in current natural gas storage aquifers [39], and it may serve effectively as a cushion gas. Nitrogen was selected because of its common use in laboratory experiments, and because it is a potential cushion gas in new storage sites [40]. Using X-ray computed tomography (XCT), we captured pore-scale fluid distribution within fractures during steady-state drainage. The recorded pressure fluctuations are analyzed to identify frequency of flow intermittency. By comparing the relative permeability and flow behavior of hydrogen, methane, and nitrogen, this study contributes to a deeper understanding of gas-brine interactions in fractured systems and offers critical data to improve the accuracy of field-scale UHS models.
@@ 2.1 Rocks and fluids
A partially karstified limestone from the Viséan (Dinantian) age was drilled from well DZH24 at the underground natural gas storage facility in Loenhout, Belgium. This limestone belongs to the Carboniferous group in the Campine-Brabant Basin of northern Belgium. The rock matrix exhibits very low porosity (0.01) and permeability (1 mD), indicating that fluid flow predominantly occurs through its fracture network. Previous reservoir characterization has shown that joints and partially open veins contribute to the overall permeability [41]. Since naturally fractured samples were unavailable for this study, artificially fractured samples were used as the best alternative for conducting our experiments.
@@
A cylindrical rock core, measuring in diameter and in length, was drilled for the experiments in this study. The core was wrapped in Teflon tape and fractured using the Brazilian tensile stress method at Ghent University’s concrete lab. This technique was selected to generate a fracture network that closely mimics natural conditions, with variable fracture width and surface roughness as shown in Figure 1. The dry fracture image and the porosity distribution along the core are shown in Figure 4(b), with the average measured porosity being 0.011. Flow through rough fractures is governed by variable aperture and relative roughness, characterized by the roughness index, , where represents the standard deviation of the aperture distribution, and is the mean aperture. Aperture was obtained from the segmented dry scan using the ‘PoreSpy’ Python toolkit [42], which calculates local thickness as shown in Figure 2(a). The aperture distribution is shown in Figure 2(b), with a mean aperture of . The relative roughness index was calculated to be 0.61.
@@
The wetting phase was prepared by dissolving 25 wt potassium iodide (Sigma Aldrich) in deionized water, which enhanced X-ray contrast for imaging purposes. The brine was selected to provide optimal contrast for X-ray imaging. While its composition differs from the natural reservoir brine, key properties influencing relative permeability, such as surface tension and viscosity, remain comparable and should not significantly impact the results. As non-wetting phases, high-purity (99.99) H2, CH4 and N2 gases (supplied by Air Liquide) were used. The thermophysical properties of these gases are listed in Table LABEL:thermophysical+properties and parameter space used in this study are shown in Figure 4(a). We maintained the same flow rate instead of matching the capillary number, as the reservoir properties are fixed. Given that nitrogen and methane are used as cushion gases, it is more practical to keep the injection rate consistent rather than focusing on the capillary number. Notably, the capillary number for all flow rates and gases is less than 10-5, indicating that the flow is capillary-dominated and exhibits similar flow regimes. It should be noted that the experimental temperature does not match the reservoir temperature due to technical limitations. However, both hydrogen and methane remain in the gaseous state, and nitrogen remains in the supercritical state under the conditions used in this study, as they would at full reservoir conditions. Consequently, the findings of this study can be reasonably extended to actual reservoir conditions.
@@ 2.2 Experimental setup and procedure
The experimental setup is shown in figure 3. Brine and gas are co-injected at increasing gas fractional flows to measure the drainage relative permeability, while keeping the constant total flow rate of 0.1 ml min-1. Brine (wetting phase) is delivered using dual syringe Brine pump (Vindum Engineering VP12K). Gas (nonwetting phase) is injected using a syringe Gas pump (ISCO 500D pump). Due to highly flammable nature of hydrogen and methane, the gas pump was stored inside a low-oxygen ( 4 ) safety cabinet. The detailed experimental procedure is as follows:
@@
The confining pressure of 15 bar was applied around the sample with deionized water using a high pressure reciprocal Confining pump (Teledyne ISCO Reaxus, USA) to secure the sample in position.
@@
The flow system was pressurized with KI brine (25 wt) upto 100 bar while maintaining 15 bar difference between inlet and confining fluid pressures
@@
Absolute permeability: The core was flooded with brine at several constant flow rates ranging from 0.1 ml min-1 to 1 ml min-1. Absolute permeability was then determined using Darcy’s law as described in equation 1. The pressure gradient was monitored at steady-state using a Keller-33PDX pressure transducer. Given the extended duration of the experiments—spanning several days—and the potential for slight alterations in fracture morphology under high-pressure flow conditions, we measured the absolute permeability prior to each set of experiments with different gases. The measured permeability values were 8.5 D for hydrogen, 7.7 D for methane, and 3.5 D for nitrogen.
@@
The fractional flow of gas is increased from 0.1 to 1 in stepwise manner. For each fractional flow, sample is scanned after steady-state is reached. The steady-state is confirmed by constant differential pressure across the length of rock sample. Fractional flow of gas is defined by equation 2. The capillary number, for each fractional flow is defined as equation 3 [46] and is given in table LABEL:Ca for each experiment in this study.
@@
Relative permeability: Relative permeability is measured at each fractional flow using equation 4.
@@ 2.3 Imaging and image analysis
X-ray imaging was performed with the ‘High Energy micro-CT Optimized for Research’ scanner (HECTOR) [47] at the center for X-ray tomography at Ghent University (UGCT). The X-ray energy and power were 160 kV and 15 W, respectively. We acquired 1301 projections with 600 ms exposure time for each projection. For each fractional flow at steady state entire sample length is scanned with imaging field of view (FOV) 2000 x 2400 x 1400 voxels. The voxel size was for each image. The Octopus software (Tescan-XRE, Belgium) was used to reconstruct the 3D image from acquired projections.
@@ 3.1 Relative permeability
The relative permeability curves were derived directly from stabilized differential pressure and saturation data, as shown in Figure 6. During drainage, as the fractional flow of gas () increases and the capillary number decreases, the system moves to lower water saturations, resulting in better-connected gas-phase displacement. This leads to an increase in gas relative permeability () and a decrease in brine relative permeability (). Previous studies report that hydrogen-water systems are water-wet with respect to calcite [48]. In such water-wet systems, gas preferentially invades regions with lower capillary entry pressures during drainage, corresponding to fracture spaces with larger apertures [49, 25, 50]. Figure 5 illustrates gas-brine occupancy at for hydrogen, methane, and nitrogen, where gas is observed occupying largest aperture spaces, while brine prefers to remain in narrower regions. This confirms that this system is water-wet. It it consistent with previous observations in porous media, where gas invades larger pores and brine remains in narrower throats [51]. Similar trends were observed at other fractional flows, as shown in the Supplementary Information S1.
@@
The relative permeability-saturation curves for the hydrogen-brine and methane-brine systems show closely aligned trends and similar endpoint saturations. Both gases exhibit low gas relative permeability (), indicating substantial phase interference caused by the heterogeneous fracture geometry (roughness) and the low viscosity ratios of the gas-brine systems (0.01 for hydrogen-brine and 0.016 for methane-brine). This behaviour suggests the occurrence of intermittent flow and snap-off events, further discussed in Section 3.3. The wetting-phase permeability () consistently remains much higher than the gas permeability, indicating that brine stays well-connected across all fractional flows. This strong water-wet condition ensures brine continuity throughout the fracture network, even as gas saturation increases. At higher water saturation (lower fractional flows, ), the gas relative permeability () is slightly higher for the methane-brine system compared to the hydrogen-brine system. This is attributed to the higher capillary number () at high wetting phase saturation, where viscous forces contribute more to the flow dynamics. The slightly higher viscosity ratio of the methane-brine system (0.016 versus 0.01 for hydrogen-brine) facilitates better connectivity in the methane front, resulting in higher gas permeability [46]. However, as the water saturation decreases (e.g., and ), capillary forces dominate, leading the values for hydrogen and methane to converge at similar endpoint saturations for hydrogen and for methane. These slight differences in relative permeability and saturation at endpoints may fall within the experimental margin of error. Notably, the saturation change from to is identical for both gases, decreasing from 0.91 to 0.6. It should be noted that the entire hydrogen-brine relative permeability experiment was repeated to assess the uncertainty in the measurements and will be discussed in section 3.4.

### v2

# Pore-scale imaging of hydrogen and methane storage in fractured aquifer rock: the impact of gas type on relative permeability

@@ 1 Introduction
The hydrogen economy is expected to play a critical role in the global transition to cleaner energy and achieving net-zero carbon emissions. A key technology supporting large-scale energy storage is Underground Hydrogen Storage (UHS). UHS enables the storage of TWh-scale energy by injecting hydrogen into depleted gas reservoirs, saline aquifers, and existing gas storage facilities. The stored (green) hydrogen can be produced by electrolysis during periods of surplus electricity generation [Tarkowski2019, Tarhan2021, Zivar2021, Thiyagarajan2022, Al-Shafi2023, Perera2023, Ali2025]. As renewable energy sources like wind and solar continue to expand [Acar2014], UHS provides a potential solution to balance renewable energy supply and demand, enhancing energy security, and helping to decarbonize energy systems. Within the broader hydrogen economy, emerging subsurface strategies such as in situ hydrogen generation from hydrocarbons [Hassanpouryouzband2025] and the conceptual framework of engineered ‘hydrogen farming’ [Hassanpouryouzband2024] further reinforces hydrogen’s role as a critical energy vector for net-zero transitions.
@@
Converting pre-existing natural gas storage sites to hydrogen storage potentially offers advantages to using depleted gas fields or saline aquifers that have not been previously deployed for storage purposes. This includes leveraging expensive site characterization campaigns, tried-and-tested reservoir and caprock behaviour under cyclical injection-retraction operations, and existing connections to gas treatment and pipeline infrastructure. Among operational and planned natural gas storage sites in Europe, 86% are depleted gas fields and 14% are aquifers, together capable of providing up to 750 TWh of hydrogen storage [HyUSPre2022]. Conversion of these storage resources would be sufficient for a mid-range scenario of 2,500 TWh of annual hydrogen demand by 2050, at a 30% storage capacity. In this paper, we focus on a notable example of such a natural gas storage site under investigation for hydrogen conversion: the Loenhout natural gas storage facility in northern Belgium. Loenhout is uniquely well-connected in Western Europe as it is part of a pipeline network linking Belgium, France, The Netherlands, Germany, Luxemburg and the United Kingdom, including international shipping terminals in the North Sea ports of Zeebrugge and Dunkirk. If converted succesfully, it would provide 2.2 TWh of storage capacity. The site comprises a partially karstified limestone aquifer from Visean age, characterized by a fracture-dominated flow system with low primary porosity [Dreesen1987, Lagrou2012, Vandenberghe2014]. Its conversion is currently under investigation in the BE-HyStore project funded by the Belgian federal Energy Transition Fund and the site operator Fluxys Belgium.
@@
Repurposing an aquifer natural gas storage site for large-scale hydrogen injection requires evaluating whether the flow of the stored gas in the pores or fractures of the storage reservoir would be affected, as this may impact injectivity and operational efficiency [Heinemann2021a, Lysyy2021]. While significant progress has been made in understanding hydrogen flow in porous media, substantial knowledge gaps remain regarding its behavior in fractured rock. This is not only important for fractured storage reservoirs such as in Loenhout, but also to assess leakage through caprock if fractures are present or generated during storage [Aftab2022, Aslannezhad2023]. A primary uncertainty in the current scientific understanding lies in hydrogen’s relative permeability compared to methane’s. Relative permeability is a key parameter for upscaled multiphase flow modeling, which quantifies how the flow of gas is impacted by the presence of (saline) groundwater in the reservoir, and vice versa [Blunt2017].
@@
Over the last two decades, several studies have shown that the relative permeability behaviour of fractured media differs significantly from that of porous materials due to its dependence on complex fluid dynamics [Phillips2020]. Nevertheless, it is currently unclear how such complexities affect fracture relative permeabilities for hydrogen compared to other gases, particularly at elevated pressures and temperatures typical for UHS. Experiments with air-water in transparent fractures have revealed strong phase interference and reduced relative permeability at intermediate saturations in rough fractures [Pruess1990, Persoff1995]. Relative permeability in such systems is highly sensitive to fracture surface roughness and aperture heterogeneity, influencing the meandering or twisting of flow paths [Chen2006, Huo2016, Nicholl2000, Yang2013, Ye2015]. Capillary-dominated flow, especially in low viscosity-ratio fluid systems such as hydrogen-brine and methane-brine, is characterized by fluid pathways that are intermittently connected, caused by snap-off events and Haines jumps [Phillips2023]. State-of-the-art fracture relative permeability models aim to capture the greater phase interference, the absence of percolating pathways, and altered flow patterns compared to porous media [Hatami2022, Watanabe2015, Wang2018, Gong2021, Eliebid2024]. The current need for understanding how the findings above translate to UHS is underscored by a lack of experimental studies on fractured rock. This is despite significant recent efforts on the corresponding behaviour of porous reservoir rocks [Yekta2018, Lysyy2022, Boon2022, Jangda2023, Higgs2024, Higgs2023, Bo2023, Rezaei2022], which are difficult to translate directly to fracture systems. Comparative studies of two-phase flow behavior of relevant fractured samples under consistent reservoir conditions are thus critical to address these uncertainties.
@@
Assessing hydrogen flow in fractured aquifers such as Loenhout is essential to evaluate its feasibility as a storage medium in comparison to methane. Therefore, our study presents the first direct measurements of drainage relative permeability for hydrogen-brine, methane-brine, and nitrogen-brine systems in fractured limestone rock at subsurface pressure conditions of 10 MPa. The investigated rock used was collected from a depth of at the Loenhout aquifer. Methane was selected because it is a good model for the stored gas in current natural gas storage aquifers [Hashemi2021], and it may serve effectively as a cushion gas. Nitrogen was selected because of its common use in laboratory experiments, and because it is a potential cushion gas in new storage sites [Zamehrian2022]. Using X-ray computed tomography (XCT), we captured pore-scale fluid distribution within fractures during steady-state drainage. The recorded pressure fluctuations are analyzed to identify frequency of flow intermittency. By comparing the relative permeability and flow behavior of hydrogen, methane, and nitrogen, this study contributes to a deeper understanding of gas-brine interactions in fractured systems and offers critical data to improve the accuracy of field-scale UHS models.
@@ 2.1 Rocks and fluids
A partially karstified limestone from the Viséan (Dinantian) age was drilled from well DZH24 at the underground natural gas storage facility in Loenhout, Belgium. This limestone belongs to the Carboniferous group in the Campine-Brabant Basin of northern Belgium. The rock matrix exhibits very low porosity (0.01) and permeability (1 mD), indicating that fluid flow predominantly occurs through its fracture network. Previous reservoir characterization has shown that joints and partially open veins contribute to the overall permeability [VanderVoet2020]. Since naturally fractured samples were unavailable for this study, artificially fractured samples were used as the best alternative for conducting our experiments.
@@
A cylindrical rock core, measuring in diameter and in length, was drilled for the experiments in this study. The core was wrapped in Teflon tape and fractured using the Brazilian tensile stress method at Ghent University’s concrete lab. This technique was selected to generate a fracture network that closely mimics natural conditions, with variable fracture width and surface roughness as shown in Figure 1(a). The dry fracture image and the porosity distribution along the core are shown in Figure 1(b), with the average measured porosity being 0.011. Flow through rough fractures is governed by variable aperture and relative roughness, characterized by the roughness index, , where represents the standard deviation of the aperture distribution, and is the mean aperture. Aperture was obtained from the segmented dry scan using the ‘PoreSpy’ Python toolkit [Gostick2019], which calculates local thickness as shown in Figure 1(c). The aperture distribution is shown in Figure 1(d), with a mean aperture of . The relative roughness index was calculated to be 0.61.
The wetting phase was prepared by dissolving 25 wt potassium iodide (Sigma Aldrich) in deionized water, which enhanced X-ray contrast for imaging purposes. The brine was selected to provide optimal contrast for X-ray imaging. The composition of brine defers from reservoir brine but sample is strongly water-wet which is also the case with realistic reservoir brine [Ringrose2020, Worden2024]. The salinity of 25 wt KI solution in terms of mole fraction is 0.0349 (34900 ppm) which falls within the typical range reported for saline aquifers in the literature [Aftab2022]. The brine was not pre-equilibrated with gas prior to injection, due to limitation of our experiment to do this for large volumes of fluids needed for co-injection. This means that the co-injected fluids equilibrate as they flow through the core. While this is relevant to near wellbore regions, the measurements may not fully representative of hydrogen-brine flow deeper in the reservoir. For example, gas dissolution into the brine could potentially lead to reduced gas connectivity close to sample inlet. However, given hydrogen’s low solubility in water and employed flow rates, the flow behavior is expected to be dominated by hydrodynamic processes, with limited dissolution.
@@
As non-wetting phases, high-purity (99.99) H2, CH4 and N2 gases (supplied by Air Liquide) were used. The thermophysical properties of these gases are listed in Table LABEL:thermophysical+properties. We maintained the same flow rate instead of matching the capillary number, as the reservoir properties are fixed. Given that nitrogen and methane are used as cushion gases, it is more practical to keep the injection rate consistent rather than focusing on the capillary number. Notably, the capillary number for all flow rates and gases is less than 10-5, indicating that the flow is capillary-dominated and exhibits similar flow regimes. It should be noted that the experimental temperature does not match the reservoir temperature due to technical limitations. However, both hydrogen and methane remain in the gaseous state, and nitrogen remains in the supercritical state under the conditions used in this study, as they would at full reservoir conditions. The interfacial tension (IFT) of the H2-water system and the viscosity of H2 gas at ambient temperature are only marginally higher by less than 6% [Hosseini2022, Chow2018] and 5% [H2tool], respectively than at reservoir temperature. Therefore, the impact of temperature on IFT and viscosity is minimal. This negligible influence is also supported by similar core-flood experiments conducted under comparable conditions [Bijay2025].
| Fluid | Density | Viscosity | Viscosity ratio | Interfacial tension |
| kg m-3 | Pa s | N m-1 | ||
| H2 [Hosseini2022] | 7 | 9.8 x 10-6 | 0.01 | 0.072 |
| CH4 [Bjorkvik2023] | 61 | 1.44 x 10-5 | 0.016 | 0.05 |
| N2 [Chow2016] | 112 | 1.5 x 10-4 | 0.16 | 0.067 |
@@ 2.2 Experimental setup and procedure
The experimental setup is shown in figure 2. Brine and gas are co-injected at increasing gas fractional flows to measure the drainage relative permeability, while keeping the constant total flow rate of 0.1 ml min-1. Brine (wetting phase) is delivered using dual syringe Brine pump (Vindum Engineering VP12K). Gas (nonwetting phase) is injected using a syringe Gas pump (ISCO 500D pump). To minimize capillary end effects, glass bead packs were placed at both the inlet and outlet of the core as fluid distributors. Due to highly flammable nature of hydrogen and methane, the gas pump was stored inside a low-oxygen (< 4 ) safety cabinet. The detailed experimental procedure is as follows:
@@
The confining pressure of 1.5 MPa was applied around the sample with deionized water using a high pressure reciprocal Confining pump (Teledyne ISCO Reaxus, USA) to secure the sample in position.
@@
The flow system was pressurized with KI brine (25 wt) upto 10 MPa while maintaining 15 bar difference between inlet and confining fluid pressures
@@
Absolute permeability: We report liquid permeability as a reference for absolute permeability, rather than gas-phase values, to avoid the influence of gas slippage. Gas permeability can vary with gas type due to the Klinkenberg effect, leading to overestimation. As shown by Klinkenberg1941 and Letham2016, liquid permeability provides a more accurate estimate of the absolute permeability. The core was flooded with brine at several constant flow rates ranging from 0.1 ml min-1 to 1 ml min-1. Absolute permeability was then determined using Darcy’s law as described in equation 1. The pressure gradient was monitored at steady-state using a Keller-33PDX pressure transducer. Given the extended duration of the experiments spanning several days and the potential for slight alterations in fracture morphology under high-pressure flow conditions, we measured the absolute permeability prior to each set of experiments with different gases and used it for obtaining relative permeability. The measured permeability values were 23 D for hydrogen - Run1 and 8.5 D for hydrogen - Run2 , 7.7 D for methane, and 3.5 D for nitrogen. This protocol ensures that relative permeability calculations for each gas system are normalized against its corresponding fracture geometry which minimizes the effect of fracture evolution. While absolute permeability variations between gas systems indicate some morphological changes occurred between experiments, the static boundary condition assumption remains valid within each individual relative permeability experiment. Furthermore, such changes in absolute permeability have been shown to have minimal impact on the shape of relative permeability curves [Eliebid2024].
@@
Figure 1(d) shows the fracture aperture distribution prior to the experiments. We did not observe any visible change in the fracture geometry in the images after the experiments. However, the measured permeability did vary, likely due to subtle pore-scale alterations that lead to fine migration during flow. Since permeability is a macroscopic property, even minor rearrangements at the pore-scale can significantly influence its value. Moreover, the image resolution is insufficient to capture voxel-scale changes in fracture geometry.
@@
The fractional flow of gas is increased from 0.1 to 1 in stepwise manner. For each fractional flow, sample is scanned after steady-state is reached. The steady-state is confirmed by constant differential pressure across the length of rock sample. Fractional flow of gas is defined by equation 2. The capillary number, for each fractional flow is defined as equation 3 [Spurin2019a] and is given in table LABEL:Ca for each experiment in this study.
@@
Relative permeability: Relative permeability is measured at each fractional flow using equation 4 [Jackson2018].
@@ 2.3 Imaging and image analysis
X-ray imaging was performed with the ‘High Energy micro-CT Optimized for Research’ scanner (HECTOR) [Masschaele2013] at the center for X-ray tomography at Ghent University (UGCT). The X-ray energy and power were 160 kV and 15 W, respectively. We acquired 1301 projections with 600 ms exposure time for each projection. For each fractional flow at steady state entire sample length is scanned with imaging field of view (FOV) 2000 x 2400 x 1400 voxels. The voxel size was for each image. The Octopus software (Tescan-XRE, Belgium) was used to reconstruct the 3D image from acquired projections.
@@ 3.1 Relative permeability
The relative permeability curves were derived directly from stabilized differential pressure and saturation data, as shown in Figure 4. During drainage, as the fractional flow of gas () increases and the capillary number decreases, the system moves to lower water saturations, resulting in better-connected gas-phase displacement. This leads to an increase in gas relative permeability () and a decrease in brine relative permeability (). Previous studies report that hydrogen-water systems are water-wet with respect to calcite [Aghaei2023]. In such water-wet systems, gas preferentially invades regions with lower capillary entry pressures during drainage, corresponding to fracture spaces with larger apertures [Karpyn2007, Phillips2023, Chen2017]. Figure 3 illustrates gas-brine occupancy at for hydrogen, methane, and nitrogen, where gas is observed occupying largest aperture spaces, while brine prefers to remain in narrower regions. This confirms that this system is water-wet. It it consistent with previous observations in porous media, where gas invades larger pores and brine remains in narrower throats [Wang2023a]. Similar trends were observed at other fractional flows, as shown in the Supplementary Information S1.
@@
The relative permeability-saturation curves for the hydrogen-brine and methane-brine systems show closely aligned trends and similar endpoint saturations. Both gases exhibit low gas relative permeability (), indicating substantial phase interference caused by the heterogeneous fracture geometry (roughness) and the low viscosity ratios of the gas-brine systems (0.01 for hydrogen-brine and 0.016 for methane-brine). This behaviour suggests the occurrence of intermittent flow and snap-off events, further discussed in Section 3.3. The wetting-phase permeability () consistently remains much higher than the gas permeability, indicating that brine stays well-connected across all fractional flows. This strong water-wet condition ensures brine continuity throughout the fracture network, even as gas saturation increases. At higher water saturation (lower fractional flows, ), the gas relative permeability () is slightly higher for the methane-brine system compared to the hydrogen-brine system. This is attributed to the higher capillary number (Cat) at high wetting phase saturation, where viscous forces contribute more to the flow dynamics. The slightly higher viscosity ratio of the methane-brine system (0.016 versus 0.01 for hydrogen-brine) facilitates better connectivity in the methane front, resulting in higher gas permeability [Spurin2019a]. However, as the water saturation decreases (e.g., and ), capillary forces dominate, leading the values for hydrogen and methane to converge at similar endpoint saturations for hydrogen and for methane. These slight differences in relative permeability and saturation at endpoints may fall within the experimental margin of error. Notably, the saturation change from to is identical for both gases, decreasing from 0.91 to 0.6. It should be noted that the entire hydrogen-brine relative permeability experiment was repeated to assess the uncertainty in the measurements and will be discussed in 3.4.